# Conformal Lyapunov Optimization: Optimal Resource Allocation under Deterministic Reliability Constraints

### v1

@@ Abstract
This paper introduces conformal Lyapunov optimization (CLO), a novel resource allocation framework for networked systems that optimizes average long-term objectives, while satisfying deterministic long-term reliability constraints. Unlike traditional Lyapunov optimization (LO), which addresses resource allocation tasks under average long-term constraints, CLO provides formal worst-case deterministic reliability guarantees. This is achieved by integrating the standard LO optimization framework with online conformal risk control (O-CRC), an adaptive update mechanism controlling long-term risks. The effectiveness of CLO is verified via experiments for hierarchal edge inference targeting image segmentation tasks in a networked computing architecture. Specifically, simulation results confirm that CLO can control reliability constraints, measured via the false negative rate of all the segmentation decisions made in the network, while at the same time minimizing the weighted sum of energy consumption and imprecision, with the latter accounting for the rate of false positives.
@@ I-A Context and Motivation
Figure 1: Illustration of the class of networks under study in which edge devices task the servers at the edge or cloud to carry out some inference tasks. Going further up towards the cloud entails larger latency and energy consumption but also a potentially higher inference quality.
In this context, it is useful to revisit existing resource allocation paradigms to assess their capability to provide optimization strategies that efficiently and reliably manage both transmission and computational resources [7]. The general goal is that of minimizing operational costs – e.g., latency, energy consumption – while ensuring strict compliance with all required service constraints.
A standard design methodology leverages Lyapunov optimization (LO) [8], a stochastic optimization tool based on queueing theory, which addresses dynamic resource allocation in networked systems. LO has been successfully applied in various contexts, including edge intelligence (EI) scenarios [9, 10]. The key advantage of LO lies in its ability to design low-complexity resource allocation procedures that minimize average network costs under long-term average constraints.
However, in applications with strict reliability requirements, ensuring average performance levels is insufficient. In fact, in such settings, the network may be required to offer strict deterministic reliability guarantees that hold even under worst-case conditions. For example, in an autonomous driving application, it may not be enough to ensure that, on average, an image classifier returns accurate predictions of street signs. Rather, it is important that the classifier outputs reliable decisions in every session. In such cases, employing the traditional LO framework may either fail to meet the required constraints or necessitate excessive complexity in the optimization process [8].
This paper proposes an extension of LO to incorporate also worst-case deterministic constraints reliability constraints. The proposed conformal Lyapunov optimization (CLO) integrates LO with online conformal risk control (O-CRC) [11, 12, 13]. O-CRC is a recently developed adaptive mechanism designed to control long-term reliability metrics in online learning environments [12]. O-CRC builds upon the conformal prediction (CP) framework [14, 11], and it is applicable to scenarios where the AI decisions take the form of a prediction set. This is the case not only of classification and regression problems, with point decisions augmented by error bars (see Figure 1 for an illustration), but also in tasks such as image segmentation or multi-label classification [11]. Specific applications include question-and-answer use cases of large language models [15, 16]. CLO endows LO with the capacity to offer deterministic performance guarantees, while extending O-CRC to address online optimization problems.
@@ I-B Related Work
From a resource optimization perspective, LO has also been employed to support goal-oriented communications, a paradigm aimed at minimizing transmission resource usage by transmitting only the essential information required to complete an inference task [27]. The work in [28] presents a general LO framework for edge-assisted goal-oriented communications, while [29] propose an LO-based resource allocation strategy leveraging convolutional neural networks. Furthermore, reference [30] explores LO-based strategies for goal-oriented neural network splitting [31].
@@
Despite the significant contributions of these works in optimizing resource allocation across various domains, none of them explicitly address optimal resource allocation under strict long-term deterministic constraints.
Conformal Prediction and Conformal Risk Control: Recent literature has highlighted the effectiveness of CP for networking applications. In [34], CP techniques – both online and offline – are applied to AI models designed to assist communication tasks such as symbol demodulation and channel estimation, while reference [35] explores the use of CP techniques for dynamic scheduling of URLLC traffic, ensuring reliability in latency-sensitive applications. In the context of spectrum access, the work [36] introduces a CRC approach for detecting occupied subbands in unlicensed spectrum access. In this work, O-CRC ensures reliable spectrum sensing by enforcing constraints on the false negative rate, thereby minimizing the likelihood of erroneously identifying an occupied spectrum portion as free.
For edge-inference scenarios, reference [37] proposes a CP-based protocol to quantify uncertainty in federated inference tasks under noisy communication channels. In a related work, the paper [38] presents a framework aimed at maximizing inference accuracy while satisfying long-term reliability and communication constraints in sensor networks equipped with a fusion center.
Among these works, only the paper [38] considers system cost optimization, while the others focus solely on satisfying long-term constraints. However, the work [38] focuses on a specific decentralized inference setting, thus not addressing the general problem of resource allocation in multi-hop edge computing networks studied in this work. Furthermore, the framework in [38] builds on online convex optimization, while the present contribution leverages LO for optimal resource allocation.
@@ I-C Main Contributions
  • We apply the framework to an edge-assisted inference scenario, where multiple edge devices perform a segmentation inference task under strict per-instance reliability constraints (see Figure 1). Our simulation results demonstrate:
@@
    • the trade-offs between average resource optimization – as provided by LO – and the satisfaction of deterministic reliability constraints – as ensured by O-CRC; and
    • the impact of enforcing deterministic reliability constraints on the trade-off between energy consumption and inference accuracy, in comparison to traditional LO-based solutions.
@@ II-A Network Model
As depicted in Figure 1, we consider a network described as a general directed graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$, with $\mathcal{N}$ denoting the set of the nodes and $\mathcal{E}\subseteq\{(n,m):n,m\in\mathcal{N},\text{with }n\neq m\}$ denoting the set of links. The set of the nodes is partitioned as
@@
where $\mathcal{U}$ denotes the set of the edge devices (ED), or users, and $\mathcal{S}$ denotes the set of the edge or cloud servers. We consider a remote inference setting scenario, where the EDs task the network with an inference problem, such as image classification or question answering under reliability constraints.
Each server in the set $\mathcal{S}$ is equipped with an inference model, such as a deep neural network or a large language model, to produce decisions on data units (DU) generated by the ED. Inference models can operate at different points on the trade-off curve between accuracy and computational cost. In particular, while we allow for a generic distribution of computational resources across servers, in practice servers can be organized in a hierarchical topology with more powerful servers being further from the ED (see Figure 1) [39].
@@ II-B Data Acquisition and Processing
We consider a discrete time axis with time-slots indexed by $t=1,2,\dots$, and each time-slot characterized by a fixed duration $\Delta T$. For each time-slot, each $k$-th ED may generate a new inference task independently with a probability $\lambda^{k}\in[0,1]$, and we denote as $A^{k}(t)\sim$ Bern($\lambda^{k}$) the random variable describing the generation of an inference task at device $k$ at time $t$. We collect the arrival processes of all the users in a random vector $\mathbf{A}(t)=\{A^{k}(t)\}_{k=1}^{K}$. We also write as $\tau^{k}(t)$ the specific inference task, e.g., an image to classify or a query to answer. In order to forward the inference task to the network, the ED produces a DU with $W^{k}$ bits encoding the task $\tau^{k}(t)$. The tasks generated at time $t$ by all the users are collected in the vector $\mathbf{T}(t)=\{\tau^{k}(t)\}_{k=1}^{K}$.
The DU encoding task $\tau^{k}(t)$ is routed to a server $s\in\mathcal{S}$, which implements the inference task. The decision is made at some later time, described by the variable $T_{\mathrm{dec}}^{k}(t)\geq t$, after the received DU is processed by server $s$. The quality of this decision depends on the complexity of the model deployed at server $s$ and on the difficulty of the task $\tau^{k}(t)$. To account for this, the loss accrued by an inference decision at server $s$ for some task $\tau$ is given by a function $L_{s}(\tau,\theta)$, which can be further controlled by a hyperparameter $\theta$.
@@ II-C Timeline
The time slots are partitioned in frames $f=0,1,\dots,$, each one composed of $S$ time slots. Thus, considering a time horizon of $T$ slots, we have $F=T/S$ frames. The frames act as monitoring time units within which the network evaluates inference performance. The rationale for defining this quantity is that, for any given application, the performance of interest is the average performance across the frame. On the basis of the average performance accrued within a time frame, future control actions may be planned. As an example, consider real-time visual tracking for micro aerial vehicles [40]. In this application, it is critical to monitor the average tracking error on suitably chosen time windows in order to take the control actions that are necessary to track the object of interest in future instants.
@@ II-D Reliability and Precision
To elaborate on the definition of the loss function $L_{s}(\tau,\theta)$, consider the image classification task depicted in Figure 1. In it, given an input image $x$, the goal of the server $s$ is to produce a subset $\mathcal{C}(x,\theta)$ of possible labels $y\in\mathcal{Y}$ as a function of the hyper-parameter $\theta$. For instance, following the conformal prediction (CP)[11] framework [41], the hyperparameter $\theta$ represents a threshold on the confidence level produced by the inference model, and the prediction set is given by
@@
$L_{s}(x,\theta)=\mathbbm{1}(y\notin\mathcal{C}(x,\theta)),$ (3)
where $\mathbbm{1}\{\cdot\}$ is the indicator function, which equals to $1$ if the argument is true and 00 otherwise. By (2), the loss (3) increases with the hyperparameter $\theta$, as required.
@@
$L_{s}(x,\theta)=\frac{\absolutevalue{y\cap\mathcal{\overline{C}}(x,\theta)}}{\absolutevalue{y}},$ (5)
where $y$ is the set of pixels including the object of interest and $\mathcal{\overline{C}}(x,\theta)$ is the complement of set $\mathcal{C}(x,\theta)$. The FNR (5) is also an increasing function of the hyperparameter $\theta$.
By the mentioned monotonicity assumption on the loss $L_{s}(\tau,\theta)$, reliability can be guaranteed by reducing the hyperparameter $\theta$. Specifically, we make the following assumption, which is satisfied in the two examples discussed above.
@@
While ensuring reliability, a small hyperparameter $\theta$ yields a less informative or precise decision. For example, in the image classification and segmentation tasks, a small $\theta$ entails larger prediction sets (2) and (4). Accordingly, there is a trade-off between reliability and precision.
To capture this trade-off, we introduce the imprecision function $F_{s}(\tau,\theta)$, which satisfies the following assumption.
@@
For example, for the classification tasks, one can adopt the imprecision measure
@@
where |$\absolutevalue{\mathcal{Y}}$ | is the size of the output space $\mathcal{Y}$, while |$\absolutevalue{\mathcal{C}(x,\theta)}$ | the size of the prediction set (2). For the segmentation problem, a relevant imprecision measure is the false positive rate (FPR)
$F_{s}(x,\theta)=\frac{\absolutevalue{\overline{y}\cap\mathcal{C}(x,\theta)}}{\absolutevalue{\overline{y}}},$ (9)
which measures the fraction of pixels assigned to the target that are actually outside the target, i.e., in the complement set $\overline{y}=\mathcal{Y}\setminus y$.
@@ II-E Transmission Model
Figure 2: Temporal evolutions of the generated DUs at ED $k$ and of the decision process at server $s$, ignoring other EDs and servers for simplicity of illustration. The time-axis at the top shows the generation process of new DUs at the ED $k$. The time-axis at the bottom shows the temporal evolution of the decision process at the server $s$. For each time instant $t$, the server may take a decision on the DU at the head of the queue $Q_{s}^{k}(t)$, whose generation time is encoded by the variable $T_{s}^{k}(t)$.
The communication model follows the standard queueing model for multi-hop wireless networks [8]. For each slot $t$, the state of each link $(n,m)\in\mathcal{E}$ is described by a non-negative variable $S_{n,m}(t)$, and the overall state matrix is $\mathbf{S}(t)=\{S_{n,m}(t)\}_{(n,m)\in\mathcal{E}}$. A power allocation matrix $\mathbf{P}(t)=\{P_{n,m}(t)\}_{(n,m)\in\mathcal{E}}$ determines the power $P_{n,m}(t)$ allocated on each edge $(n,m)\in\mathcal{E}$ at time $t$. Thus, the overall power consumption of the $n$-th node in the network is given by the sum
@@
for some capacity function $C_{n,m}(\cdot)$. For example, in the absence of interference, and transmission on AWGN channels, the capacity function can be chosen according to Shannon theory as [43]
@@
where $B_{n,m}$ represents the transmission bandwidth for the link $(m,n)$, while $N_{0}$ is the noise power spectral density.
Recalling that $W^{k}$ represents the size in bits of the DUs generated by the $k$-th user, the transmission time of a data-unit generated by the $k$-th user across the link $(n,m)$ is given by
@@
which we assume to be no longer than the duration $\Delta T$ of the time slot. Thus, the energy required to forward a DU of the $k$-th ED at the $t$-th slot is computed as the product
@@
Define as $R_{n,m}^{k}(t)$ the binary variable that indicates if the link $(n,m)$ is used for the transmission of a DU by the ED $k$ in the time slot $t$, i.e.,
$R_{n,m}^{k}(t)=\begin{cases}1,&\text{link $(n,m)$ carries a DU of the $k$-th ED}\\ 0,&\text{otherwise}.\end{cases}$ (15)
We impose the constraint on the maximum number of DUs that can be sent on any link $(n,m)$
$\sum_{k=1}^{K}R_{n,m}^{k}(t)\leq R_{\mathrm{max}}^{k}\hskip 22.0pt.$ (16)

### v5

@@ Abstract
This paper introduces conformal Lyapunov optimization (CLO), a novel resource allocation framework for networked systems that optimizes average long-term objectives, while satisfying deterministic long-term reliability constraints. Unlike traditional Lyapunov optimization (LO), which addresses resource allocation tasks under average long-term constraints, CLO provides formal worst-case deterministic reliability guarantees. This is achieved by integrating the standard LO optimization framework with online conformal risk control (O-CRC), an adaptive update mechanism controlling long-term risks. The effectiveness of CLO is verified via experiments for hierarchal edge inference targeting image segmentation tasks in a networked computing architecture. Specifically, simulation results confirm that CLO can control reliability constraints, measured via the false negative rate of all the segmentation decisions made in the network, while at the same time minimizing the weighted sum of energy consumption and precision loss, with the latter accounting for the rate of false positives.
@@ I-A Context and Motivation
Figure 1: Edge devices task the servers at the network edge, or in the cloud, to carry out some inference task. Cloud servers typically entails larger latency and energy consumption but, potentially, also a better inference.
In this context, it is useful to revisit existing resource allocation paradigms to assess their capability to provide optimization strategies that efficiently and reliably manage both transmission and computational resources [7]. The general goal is minimizing operational costs – e.g., latency, energy consumption – while ensuring strict compliance with all required service constraints.
A standard design methodology leverages Lyapunov optimization (LO) [8], a stochastic optimization tool based on queuing theory, which addresses dynamic resource allocation in networked systems. LO has been successfully applied in various contexts, including edge intelligence (EI) scenarios [9, 10]. The key advantage of LO lies in its ability to design low-complexity resource allocation procedures that minimize average network costs, under long-term average constraints.
However, in applications with strict reliability requirements, ensuring average performance levels is insufficient. In fact, in such settings, the network may be required to offer strict deterministic reliability guarantees that hold even under worst-case conditions. For example, in an autonomous driving application, it may be not enough to ensure that, on average, an image classifier returns accurate predictions of street signs. Rather, it is important that the classifier outputs reliable decisions in every session. In such cases, employing traditional LO frameworks may either fail to meet the required constraints or request an excessively complex optimization process [8].
This paper proposes an extension of LO, named conformal Lyapunov optimization (CLO), which incorporates also worst-case deterministic reliability constraints, by integrating LO with online conformal risk control (O-CRC) [11, 12, 13]. O-CRC is a recently developed adaptive mechanism designed to control long-term reliability metrics in online learning environments [12]. O-CRC builds upon the conformal prediction (CP) framework [14, 11], and it is applicable to scenarios where the AI decisions take the form of a prediction set. This is the case not only of classification and regression problems, with point decisions augmented by error bars (see Figure 1 for an illustration), but also in tasks such as image segmentation or multi-label classification [11]. Specific applications include question-and-answer use cases of large language models [15, 16]. CLO endows LO with the capacity to offer deterministic performance guarantees, while extending O-CRC to address online optimization problems.
@@ I-B Related Work
From a resource optimization perspective, LO has also been employed to support goal-oriented communications, a paradigm aimed at minimizing transmission resource usage by transmitting only the essential information required to complete an inference task [27]. The work in [28] presents a general LO framework for edge-assisted goal-oriented communications, while [29] considers an LO-based resource allocation strategy leveraging convolutional neural networks. Furthermore, reference [30] explores LO-based strategies for goal-oriented neural network splitting [31].
@@
Despite the significant contributions of these works in optimizing networked resource allocation across various domains, none of them explicitly address optimal resource allocation under strict long-term deterministic constraints.
Conformal Prediction and Conformal Risk Control: Recent literature has highlighted the effectiveness of CP for networking applications. In [34], CP techniques – both online and offline – are applied to AI models designed to assist communication tasks, such as symbol demodulation and channel estimation, while [35] explores the use of CP techniques for dynamic scheduling of URLLC traffic, ensuring reliability in latency-sensitive applications. In the context of spectrum access, authors in [36] introduce a CRC approach for detecting occupied subbands in unlicensed spectrum access. Therein, O-CRC ensures reliable spectrum sensing by enforcing constraints on the false negative rate, thereby minimizing the likelihood of erroneously identifying an occupied spectrum portion as free.
For edge-inference scenarios, [37] proposes a CP-based protocol to quantify uncertainty in federated inference tasks under noisy communication channels. In a related work, [38] presents a framework aimed at maximizing inference accuracy while satisfying long-term reliability and communication constraints in sensor networks equipped with a fusion center.
Among these works, only [38] considers system cost optimization, while the others focus solely on satisfying long-term constraints. However, [38] focuses on a specific decentralized inference setting, thus not addressing the general problem of resource allocation in multi-hop edge computing networks studied herein. Furthermore, the framework in [38] builds on online convex optimization, while the present contribution leverages LO for optimal resource allocation.
@@ I-C Main Contributions
  • We apply the framework to an edge-assisted inference scenario, where multiple devices perform their own inference task (i.e, segmentation), possibly offloading computations to (edge/cloud) servers, under strict per-instance reliability constraints (see Figure 1). The simulation results show:
@@
    • the trade-offs between average resource optimization (granted by LO), and the satisfaction of deterministic reliability constraints (ensured by O-CRC);
    • the impact of extra deterministic reliability constraints on classical LO policies, on the trade-off between energy consumption and inference accuracy.
@@ II-A Network Model
As depicted in Figure 1, we consider a network described as a directed graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$, with $\mathcal{N}$ denoting the set of the nodes and $\mathcal{E}\subseteq\{(n,m):n,m\in\mathcal{N},\text{with }n\neq m\}$ denoting the set of links. The set of the nodes is partitioned as
@@
where $\mathcal{U}$ denotes the set of the edge devices (ED), or users, and $\mathcal{S}$ denotes the set of the edge or cloud servers. We consider a remote inference setting scenario, where the EDs may decide to load the network with an inference problem, such as image classification, or question answering, under reliability constraints.
Each server in the set $\mathcal{S}$ is equipped with an inference model, such as a deep neural network or a large language model, to produce decisions on data units (DU) generated by the EDs. Inference models can operate at different points on the trade-off curve between accuracy and computational cost. In particular, while we allow for a generic distribution of computational resources across servers, in practice servers can be organized in a hierarchical topology with more powerful servers being further from the ED (see Figure 1) [39], and possibly affected by a higher (transmission) latency.
@@ II-B Data Acquisition and Processing
We consider a discrete-time axis with time-slots indexed by $t=1,2,\dots$, and each time-slot characterized by a fixed duration $\delta$. For each time-slot, each $k$-th ED may generate a new inference task $\tau^{k}(t)$, e.g., an image to classify or a query to answer, independently from each other, and with a probability $\lambda^{k}\in[0,1]$. We denote as $A^{k}(t)\in\{0,1\}$ the binary random variable indicating the arrival of a new task $\tau^{k}(t)$, and of the corresponding data-unit (DU) for the $k$-th device at $t$-th slot, and we collect the arrival processes of all the users in a random vector $\mathbf{A}(t)=\{A^{k}(t)\}_{k=1}^{K}$. In order to forward the inference task to the network, the ED produces a DU with $W^{k}$ bits encoding the task $\tau^{k}(t)$. The tasks generated at time $t$ by all the users are collected in the vector $\mathbf{T}(t)=\{\tau^{k}(t)\}_{k=1}^{K}$.
The DU encoding task $\tau^{k}(t)$ is routed to a server $s\in\mathcal{S}$, which implements the inference task. The decision is made at some later time, described by the variable $T_{\mathrm{dec}}^{k}(t)\geq t$, after the received DU is processed by server $s$. The quality of this decision depends on the complexity of the model deployed at server $s$ and on the difficulty of the task $\tau^{k}(t)$. This decision quality for any inference task $\tau$ at each server $s$, is summarized by a loss function $L_{s}(\tau,\theta)$, which is assumed to be further controllable by a hyperparameter $\theta$.
@@ II-C Timeline
The time slots are partitioned in frames $f=0,1,\dots,$, each one composed of $S$ time slots. Thus, considering a time horizon of $T$ slots, we have $F=T/S$ frames. The frames act as monitoring time units within which the network evaluates inference performance. The rationale for defining this quantity is that, for any given application, the performance of interest is the average performance across the frame. On the basis of the average performance accrued within a frame, future control actions may be planned. As an example, consider real-time visual tracking for micro aerial vehicles [40]. In this application, it is critical to monitor the average tracking error on suitably chosen time windows in order to take the control actions that are necessary to track the object of interest in future instants.
@@ II-D Reliability and Precision
To elaborate on the definition of the loss function $L_{s}(\tau,\theta)$, consider the image classification task depicted in Figure 1. In this case, given an input image $x$, the goal of the server $s$ is to produce a subset $\mathcal{C}(x,\theta)$ of possible labels $y\in\mathcal{Y}$ as a function of the hyperparameter $\theta$. For instance, following the conformal prediction (CP)[11] framework [41], the hyperparameter $\theta$ represents a threshold on the confidence level produced by the inference model, and the prediction set is given by
@@
$L_{s}(x,\theta)=\mathbbm{1}(y_{\mathrm{true}}\notin\mathcal{C}(x,\theta)),$ (3)
where $y_{\mathrm{true}}$ is the true label associated to the input, and $\mathbbm{1}\{\cdot\}$ is the indicator function, which equals to $1$ if the argument is true and 00 otherwise. By (2), the loss (3) increases with the hyperparameter $\theta$, as required.
@@
$L_{s}(x,\theta)=\frac{\absolutevalue{y_{\mathrm{true}}\cap\mathcal{\overline{C}}(x,\theta)}}{\absolutevalue{y_{\mathrm{true}}}},$ (5)
where $y_{\mathrm{true}}$ is the set of pixels including the object of interest and $\mathcal{\overline{C}}(x,\theta)$ is the complement of set $\mathcal{C}(x,\theta)$. The FNR (5) is also an increasing function of the hyperparameter $\theta$.
By the mentioned monotonicity assumption on the loss $L_{s}(\tau,\theta)$, a higher reliability (e.g., a lower loss) can be guaranteed by reducing the hyperparameter $\theta$. Specifically, we make the following assumption, which is satisfied in the two examples discussed above.
@@
While increasing reliability, a smaller hyperparameter $\theta$ yields a less informative, or precise, decision. For example, in image classification and segmentation, a small $\theta$ entails larger prediction sets (2) and (4). Accordingly, there is a trade-off between reliability (e.g., true pixels in the prediction set) and precision (e.g., correct pixels w.r.t. the set cardinality).
To capture this trade-off, we introduce the precision loss $F_{s}(\tau,\theta)$, which satisfies the following assumption.
@@
For example, for classification tasks, one can adopt the precision loss
@@
where |$\absolutevalue{\mathcal{Y}}$ | is the size of the output space $\mathcal{Y}$, while |$\absolutevalue{\mathcal{C}(x,\theta)}$ | the size of the prediction set (2). For image segmentation, a widely used precision loss is the false positive rate (FPR)
$F_{s}(x,\theta)=\frac{\absolutevalue{\overline{y}_{\mathrm{true}}\cap\mathcal{C}(x,\theta)}}{\absolutevalue{\overline{y}_{\mathrm{true}}}},$ (9)
i.e., the fraction of pixels of the estimated target that are outside the true target, e.g., in the set $\overline{y}_{\mathrm{true}}=\mathcal{Y}\setminus y_{\mathrm{true}}$.
Appendix A reports the proofs of monotonicity for the presented precision and reliability losses..
@@ II-E Transmission Model
Figure 2: Sequence of DUs generated at $k$-th ED and associated decisions at server $s$ (ignoring other EDs and servers). At any time $t$, the server may decide on the DU at the head of its queue $Q_{s}^{k}(t)$, whose generation time is encoded by $T_{s}^{k}(t)$.
The transmission phase follows a standard queuing model for multi-hop wireless networks [8]. In each slot $t$, the link $(n,m)\in\mathcal{E}$ is described by the channel state $S_{n,m}(t)$, and the overall state matrix is $\mathbf{S}(t)=\{S_{n,m}(t)\}_{(n,m)\in\mathcal{E}}$. A power allocation matrix $\mathbf{P}(t)=\{P_{n,m}(t)\}_{(n,m)\in\mathcal{E}}$ determines the power $P_{n,m}(t)$ allocated on each edge $(n,m)$ at time $t$. The overall power consumption of the $n$-th node in the network is given by the sum
@@
for some capacity function $C_{n,m}(\cdot)$. For example, in AWGN channels without interference, according to Shannon theory the capacity function can be chosen as [43]
@@
where $B_{n,m}$ represents the transmission bandwidth for the link $(n,m)$, while $N_{0}$ is the noise power spectral density.
Recalling that $W^{k}$ represents the size in bits of the DUs generated by the $k$-th user, the transmission delay of a DU generated by the $k$-th user across the link $(n,m)$ is given by
@@
which we assume to be no longer than the duration $\delta$ of the time slot. Thus, the energy required to forward a DU of the $k$-th ED at the $t$-th slot is expressed by
@@
Indicating with $R_{n,m}^{k}(t)$ the binary variable capturing if the link $(n,m)$ is used for the transmission of a DU by the $k$-th ED in the time slot $t$, i.e.,
$R_{n,m}^{k}(t)=\begin{cases}1,&\text{link $(n,m)$ carries a DU of the $k$-th ED}\\ 0,&\text{otherwise},\end{cases}$ (15)
we can constraint the maximum number of DUs that can be sent on any link $(n,m)$, by
$\sum_{k=1}^{K}R_{n,m}^{k}(t)\leq R_{n,m}^{\max}\hskip 12.0pt\forall\;(n,m),t.\hskip 22.0pt.$ (16)